OpenStudy (anonymous):

Let f(x)=lx^2-35l. Find the average rate of change of f(x) from x=7 to x=4.

OpenStudy (anonymous):

i know how to do rate of change but this prob is a little different

OpenStudy (asnaseer):

average rate of change is the change in f(x) from x1 to x2 divided by the interval size |x1-x2]

OpenStudy (anonymous):

ok let me show u what u did: f(7)=l(7)^2-35l=l49-35l=l14l=14 f(4)=l(4)^2-35l=l16-35l=l-19l=19 and then 19-14 ? ...that's all i had

OpenStudy (anonymous):

i mean what I did

OpenStudy (anonymous):

my bad

OpenStudy (asnaseer):

yes @mariomintchev, you are almost there, its:\[(14-19)/3=-5/3\]

OpenStudy (anonymous):

why isnt it 19-14

OpenStudy (asnaseer):

rate of change means "how fast is one value changing with respect to another". e.g. think of instantaneous rate of change - that would be dy/dx

OpenStudy (chriss):

\[(f(7)-f(4))/7-4\] (14-(-19))/3 33/3 11

OpenStudy (anonymous):

where did u guys get the 3 from?

OpenStudy (anonymous):

o ok nevermind!

OpenStudy (chriss):

x1-x2

OpenStudy (asnaseer):

here you are asked for an average rate of change - so it is the change in f(x) divided by the change in x @ChrisS - I think you missed the "magnitude" operator in f(x)

OpenStudy (anonymous):

its positive 19 not -19

OpenStudy (chriss):

right, I see that now.. my bad

OpenStudy (anonymous):

also the answer would be -5/4 not -5/3

OpenStudy (anonymous):

o nevermind again

OpenStudy (asnaseer):

:-)

OpenStudy (anonymous):

THANKS :)